OpenStudy (anonymous):

SA=11^(2)π+11π√(11^(2)+8^(2)

OpenStudy (anonymous):

Are you wondering how to simplify this?

OpenStudy (anonymous):

yes really how to to it on caculator

OpenStudy (anonymous):

really how to input it because spose to sa=850.17cm^2 but when i do it on caculator am not getting that answer

OpenStudy (anonymous):

\[\left( 121+11\sqrt{85} \right)\pi\]

OpenStudy (anonymous):

Your expression doesn't equal 850.17 cm^2 so maybe your expression isn't quite right.

OpenStudy (anonymous):

SA=11^(2)`+11`~(11^(2)+8^(2))

OpenStudy (anonymous):

comes to 850.17

OpenStudy (anonymous):

but i never can get that

OpenStudy (anonymous):

121*pi + 11*sqrt(185)*pi?

OpenStudy (anonymous):

Oh you're right - I entered it wrong the first time.

OpenStudy (anonymous):

This is what you should enter into your calculator: \[(121+11\sqrt{185})*\pi\]

OpenStudy (anonymous):

okay any way can brake it down how you get the answer

OpenStudy (anonymous):

how you get 121 and 185

OpenStudy (anonymous):

11*11 is 121 121 + 64 = 185

OpenStudy (anonymous):

how get 64 lol

OpenStudy (anonymous):

8*8

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how i get the .17

OpenStudy (anonymous):

When you multiply by pi you'll get extra decimal numbers.

OpenStudy (anonymous):

all okay

OpenStudy (anonymous):

sorry i am putting them in and still not getting that answer with it not entering it in right

OpenStudy (anonymous):

okay so enter it in this order: 11*sqrt(185) + 121 = ..... then take that answer and multiply it by pi (if you have that button) or 3.14 if you don't have a pi button.